# Out-of-distribution generalization via composition: a lens through induction heads in Transformers

### v1

@@ 3.2 Common bridge representation hypothesis
How do compositions work in LLMs beyond the synthetic example? With a microscopic analysis focusing on the copying task, we posit the following Common Bridge Representation (CBR) hypothesis.
@@
This latent subspace can be viewed as a sophisticated example of feature superposition [25] for compositional tasks.
We experiment on a variety of LLMs and highlight key findings in Figure 5(b)–(f), where we use GPT-2 as a recurring example and summarize results of other models. Top-scoring PTHs and IHs are distributed across different layers of LLMs, though more PTHs appear in early layers. We sample sequences of the format $({\bm{s}}^{\#},{\bm{s}}^{\#},{\bm{s}}^{\#})$ and calculate the average token-wise probability/accuracy for predicting the third segment ${\bm{s}}^{\#}$. See experiment details in Section C.
@@
For both experiments, we apply screening to top PTHs and IHs based on diagonal scores, resulting in an average of $9$ effective PTHs and $7$ effective IHs. The first intervention experiment involves shuffling heads. We randomly permute matrices ${\bm{W}}_{\mathrm{QK}}$ within the list of IHs, yielding an edited model (“shuffle within”). As comparison, we replace each ${\bm{W}}_{\mathrm{QK}}$ within the list wby a random ${\bm{W}}_{\mathrm{QK}}$ outside the list, yielding another edited model (“shuffle outside”). In a parallel experiment, we shuffle ${\bm{W}}_{\mathrm{OV}}$ circuits within PTHs similarly. We evaluate the original model and edited models by calculating the average probability of predicting correct tokens.
The second intervention experiment involves projection of weight matrices. First, we stack the QK matrices from IHs into a matrix $[{\bm{W}}_{\mathrm{QK}}^{1},\ldots,{\bm{W}}_{\mathrm{QK}}^{K}]$ and extract the top right singular vectors ${\bm{V}}\in\mathbb{R}^{d\times r}$ for a pre-specified $r$. We call the column linear span of ${\bm{V}}$ as the bridge subspace. Then, we edit $25\%$ of all attention heads by weight projection ${\bm{W}}_{\mathrm{QK}}\leftarrow{\bm{W}}_{\mathrm{QK}}{\bm{V}}{\bm{V}}^{\top}$. After the edit (‘keep’), the attention calculation can only use the component of embeddings within the bridge subspace. In a parallel experiment, we make a projection edit (‘remove’) with an orthogonal complement ${\bm{W}}_{\mathrm{QK}}\leftarrow{\bm{W}}_{\mathrm{QK}}({\bm{I}}_{d}-{\bm{V}}{\bm{V}}^{\top})$ to force attention calculation not to use component in the bridge subspace. We evaluate the edited models by calculating the average accuracy of predicting correct tokens.
@@ 4 Related work
Mechanistic interpretability (MI) aims to provide microscopic interpretability of inner workings of Transformers [57]. In particular, [26, 58] proposed IHs as crucial components of Transformers. In particular, they suggested that matching OV circuits and QK circuits is crucial for ICL. A line of research extends IHs in various aspects [84, 66, 52, 74, 27, 5, 29, 4]. Compared with the existing literature, we conducted extensive experiments on both small Transformers and a variety of LLMs, demonstrating that IHs are crucial components for many reasoning tasks. Further, we propose that subspace matching—and more broadly common bridge representation hypothesis—as the compositional mechanism.
@@
Historically, studies of generalization on novel domains focus on extrapolation, distribution shift, domain adaptation, etc. Since GPT-3 [16], recent studies of OOD generalization focus on arithmetic and algebraic tasks [43, 96], formal language and deductive reasoning [71, 67, 77, 83], learning boolean functions [1], etc. Our copying task is an example of length generalization [7, 99]. Compositional tasks are strongly related to reasoning abilities of LLMs, such as arithmetic tasks [24, 42], formal language [32], logical rules [14, 91], and so on. Despite recent insights [10, 24], a systematic analysis remains challenging. Our work is a first step toward understanding sophisticated compositional capabilities of LLMs.
@@
Linear Representation Hypothesis (LRH) states that monosemantic (meaning basic or atomic) concepts are represented by linear subspaces (or half-spaces) in embeddings [11, 25, 63]. This hypothesis is empirically verified in not only in LLMs but also in simpler statistical models [53, 64]. LRH treats subspaces as fundamental units for representing linguistic concepts, but it does explain how models solve reasoning tasks or achieve OOD generalization. Our CBR hypothesis furthers this view by linking intermediate outputs in compositions to interpretable subspaces.
@@ 5 Limitations and future work
Second, we did not develop insights to explain the emergence of the bridge subspace during training. The sharp transition in prediction accuracy is related to the emergent abilities of LLMs observed in broader contexts [86]. In simpler settings, feedforward networks learning algebraic rules also exhibit phase transitions from memorization to generalization, a phenomenon known as Grokking [65, 56, 98, 48, 46, 51]. Analyzing the training dynamics of Transformers for copying and other compositional tasks would be an interesting avenue for further research
@@ B.3 One-layer Transformer
Figure 1 shows that one-layer Transformers do not learn the rule of copying under 20K training steps. We used the same model architecture and hyperparameters for training as detailed in Section C.1.
@@ B.4 Larger Transformers
The results of the three models are similar to the minimal model (2-layer 1-head Transformer with no MLP) we analyzed in the paper. See plots in Section D.1.
@@ B.5 Details about models
There are two major variants of positional embedding: absolute positional embedding (APE) and rotary embedding (RoPE). APE is proposed in the original Transformer paper [81] and used widely in early LLMs such as GPT-2. The input vector to a Transformer is obtained by adding a token-specific embedding vector and a position-specific embedding vector, thus encoding both token information and positional information. RoPE [76] is often used in recent LLMs such as Llama-2. The positional information is not encoded at the input layer; instead, RoPE directly modifies the attention by replacing ${\bm{W}}^{q}({\bm{W}}^{k})^{\top}$ in Eq. 5 with ${\bm{W}}^{q}{\bm{R}}_{\Delta t}({\bm{W}}^{k})^{\top}$ where ${\bm{R}}_{\Delta t}\in\mathbb{R}^{d_{\mathrm{head}}\times d_{\mathrm{head}}}$ is a matrix that depends on relative distance of two tokens. In particular, if $\Delta t=0$ then ${\bm{R}}_{\Delta t}={\bm{I}}_{d}$. Our analysis in Section 2 and 3 essentially treats $\Delta t=0$.
Section D.5 explores the impact of ${\bm{R}}_{\Delta t}$ on our results. We did not find significant differences when we repeat our measurements with a nonzero $\Delta t$.
@@ B.6 Details about measurements
Appendix C Experiment details for experiments with LLMs
C.1 Models
@@
  • Llama2-70B: 64 layers, 80 attention heads, and a hidden size of 8192 (for chain-of-thought experiments in Section E.1).
@@
C.2 Prompt examples
Here we list a few examples for the three tasks in Section 3.1.
Fuzzy copying.
Prompt   : bear snake fox poppy plate butterfly caterpillar boy
@@ C.2 Prompt examples
Solution : AQUARIUM_FISH MOTORCYCLE
Prompt   : willow_tree bee orchid man plain bicycle pear crab otter possum
WILLOW_TREE BEE ORCHID MAN PLAIN BICYCLE PEAR CRAB
Solution : OTTER POSSUM
Indirect Object Identification.
@@
Prompt   : Then, Anna and Matthew went to the station. Anna gave a basketball to
Solution : Matthew
@@
Prompt   : Then, &^ and #\$ went to the station. &^ gave a basketball to
Solution : #\$
In-context Learning.
@@
Prompt   : baseball is sport, celery is plant, sheep is animal,
@@
Prompt   : baseball is \$#, celery is !%, sheep is &*, volleyball is \$#,
@@
Solution : !%
C.3 Shuffle intervention experiment
Evaluation.
$p_{2L+t}\big{(}s_{t}^{\#}\big{|}({\bm{s}}^{\#},{\bm{s}}^{\#},{\bm{s}}_{<t}\big{)}.$ (7)
For GPT2, we gather the above probability for $t\in\{6,7,8,\ldots,L\}$ and every input sequence. Finally, in the left plot in Figure 5(e), we compare the histogram of the prediction probabilities of the original model, the edited model with shuffling, and the edited random baseline.
C.4 Projection intervention experiment
Calculating bridge subspace.
As mentioned in Section 3.2, we calculated the bridge subspace of rank $r$ based on the top-$K$ ${\bm{W}}_{\mathrm{QK}}$ from the list of IHs. We chose to edit $25\%$ of all heads using the projection matrices ${\bm{V}}{\bm{V}}^{\top}$ or ${\bm{I}}_{d}-{\bm{V}}{\bm{V}}^{\top}$ because we believe that $25\%$ of all heads are likely to be relevant to the copying task. These heads are selected as the $25\%$ top-ranking heads according to the IH attention score.
Evaluation.
50 80 150 170 200 200 200 200 200 220
Table 1: The rank $r$ of ${\bm{V}}$ used in the projection calculation.
C.5 Connection to spiked matrices
Appendix D Additional results for the synthetic example
D.1 Larger models show similar sharp transitions
We believe that the 2-layer 1-head Transformer with no MLP is a minimal yet representative model for analyzing the sharp transition. We show similar results on three larger Transformer models in Figure 7 to 10.
Figure 10: Progress measures for 2-layer 1-head Transformer with MLP.
D.2 Progress measures under varying pool sizes
We consider the same model/training settings as in Section 2 but we choose two different pool sizes $S=1000$ and $S=100$. In Figure 11 and 12, the progress measures under $S=1000$ are similar to Figure 2 but very different under $S=100$.
D.3 Attention matrices
The heatmaps in Figure 3 indicate that the memorizing model ($S=100$) are qualitatively different from that generalizing model ($S=1000$). In Figure 13 and 14, we further plot attention matrices ${\bm{A}}$ based on one OOD sequence at training step 16K to support our claim that the memorizing model does not learn compositional structure for OOD generalization.
D.4 Measurements for subspace matching
$\sigma_{\max}({\bm{U}}^{\top}{\bm{V}})$
  • Vary the rank parameter $r$.
We find that the two alternative measurement yield qualitatively similar results; see Figure 15 and 16.
D.5 Rotary positional embedding
${\bm{W}}_{\mathrm{QK},\Delta t}={\bm{W}}^{q}{\bm{R}}_{\Delta t}({\bm{W}}^{k})^{\top}$
Appendix E Additional results for the LLMs experiments
E.1 Additional results for symbolized language reasoning
Apart from the three tasks in Section 3.1, we offer an additional task. Chain-of-thought (CoT) reasoning empirically works well on mathematical tasks. It usually involves a step-by-step manner. We demonstrate that IHs also have a large impact on CoT reasoning.
  1. 4.
    Chain-of-thought (CoT): we sample instances from the GSM8K Dataset [19] which contain grade school math problems.
For instance, “Ivan had \$10 and spent 1/5 of it on cupcakes. He then spent some money on a milkshake and had only \$3 left. How much is the milkshake?” We concatenate pairs of a problem description and a step-by-step solution as in-context examples in the prompt. We use Llama2-70B as the pretrained LLM to generate answers. We measure the accuracy of the original/edited model by generating $256$ tokens, extracting the first numeric number from generated tokens, and comparing it with the correct number.
Each problem consists of several base arithmetic subproblems. While the subproblems are likely ID, due to the freedom of combining the subproblems, the problem may be OOD. Figure 21 shows that removal of IHs causes significant accuracy drop.
Figure 21: Chain-of-thought depends on IHs. Accuracy drops as we increase the number of removed IHs from the model.
E.2 Details about PTHs and IHs in LLMs

### v2

@@ 4.3 Common bridge representation hypothesis
How do compositions work in LLMs beyond the synthetic example? We consider four tasks: (i) Copying, (ii) Fuzzy Copying, (iii) IOI, and (iv) ICL. For illustrative purposes, we report results for the copying task. The results for the other three tasks are similar, which can be found in Appendix F. With a microscopic analysis, we posit the following Common Bridge Representation (CBR) hypothesis.
@@
The CBR hypothesis is an extension of linear representation hypothesis to compositional tasks. We give a mathematical statement of the CBR hypothesis in the ideal scenario. For a matrix ${\bm{W}}_{0}\in\mathbb{R}^{d\times d}$, we denote by $\mathrm{span}({\bm{W}}_{0}):=\{{\bm{W}}_{0}{\bm{x}}:{\bm{x}}\in\mathbb{R}^{d}\}$ the column span of ${\bm{W}}_{0}$. Let ${\bm{W}}_{\mathrm{OV},j},{\bm{W}}_{\mathrm{QK},k}\in\mathbb{R}^{d\times d}$ be the matrices in a task-relevant OV/QK circuit in an earlier/later layer respectively. Then, for a compositional task, there exists a low-dimensional subspace $\mathcal{V}\subset\mathbb{R}^{d}$ such that
$\mathcal{V}=\mathrm{span}({\bm{W}}_{\mathrm{OV},j})=\mathrm{span}({\bm{W}}_{\mathrm{QK},k}^{\top}).$ (7)
Intuitively, the output of an OV circuit has the form ${\bm{W}}_{\mathrm{OV},j}{\bm{x}}$, which lies in $\mathcal{V}$. It needs to be processed by a later QK circuit through ${\bm{W}}_{\mathrm{QK},k}{\bm{x}}$ to match relevant hidden states.
LLMs in practice are trained with stochastic gradient descent, so we do not expect Eqn. 7 to hold exactly. Nevertheless, it motivates us in our experiment to estimate $\mathcal{V}$ from a collection of relevant QK circuits: given $({\bm{W}}_{\mathrm{QK},k})_{k}^{K}$, we apply singular value decomposition to the stacked matrix $[{\bm{W}}_{\mathrm{QK},1},\ldots,{\bm{W}}_{\mathrm{QK},K}]$ and use the principal right singular subspace as an estimate of $\mathcal{V}$.
We experiment on a variety of LLMs and highlight key findings in Figure 7(b)–(f), where we use GPT-2 as a recurring example and summarize results of other models. Top-scoring PTHs and IHs are distributed across different layers of LLMs, though more PTHs appear in early layers. We sample sequences of the format $({\bm{s}}^{\#},{\bm{s}}^{\#},{\bm{s}}^{\#})$ and calculate the average token-wise probability/accuracy for predicting the third segment ${\bm{s}}^{\#}$. See experiment details in the Appendix.
@@
For both experiments, we apply screening to top PTHs and IHs based on diagonal scores, resulting in an average of $9$ effective PTHs and $7$ effective IHs. The first intervention experiment involves shuffling heads. We randomly permute matrices ${\bm{W}}_{\mathrm{QK}}$ within the list of IHs, yielding an edited model (“shuffle within”). As comparison, we replace each ${\bm{W}}_{\mathrm{QK}}$ within the list by a random ${\bm{W}}_{\mathrm{QK}}$ outside the list, yielding another edited model (“shuffle outside”). In a parallel experiment, we shuffle ${\bm{W}}_{\mathrm{OV}}$ circuits within PTHs similarly. We evaluate the original model and edited models by calculating the average probability of predicting correct tokens.
The second intervention experiment involves projection of weight matrices based on the discussion on Eqn. 7. First, we stack the QK matrices from IHs into a matrix $[{\bm{W}}_{\mathrm{QK}}^{1},\ldots,{\bm{W}}_{\mathrm{QK}}^{K}]$ and extract the top right singular vectors ${\bm{V}}\in\mathbb{R}^{d\times r}$ for a pre-specified $r$. We call the column linear span of ${\bm{V}}$ as the bridge subspace. Then, we edit $25\%$ of all attention heads by weight projection ${\bm{W}}_{\mathrm{QK}}\leftarrow{\bm{W}}_{\mathrm{QK}}{\bm{V}}{\bm{V}}^{\top}$. After the edit (‘keep’), the attention calculation can only use the component of embeddings within the bridge subspace. In a parallel experiment, we make a projection edit (‘remove’) with an orthogonal complement ${\bm{W}}_{\mathrm{QK}}\leftarrow{\bm{W}}_{\mathrm{QK}}({\bm{I}}_{d}-{\bm{V}}{\bm{V}}^{\top})$ to force attention calculation not to use component in the bridge subspace. We evaluate the edited models by calculating the average accuracy of predicting correct tokens.
@@ 5 Related work
Mechanistic interpretability (MI) aims to provide microscopic interpretability of inner workings of Transformers [54]. In particular, [25, 55] proposed IHs as crucial components of Transformers. In particular, they suggested that matching OV circuits and QK circuits is crucial for ICL. A line of empirical research extends IHs in various aspects [80, 61, 48, 68, 26, 4, 29, 3], and [53] provides theoretical analysis of IHs assuming a latent causal graph. Compared with the existing literature, we conducted extensive experiments on both small Transformers and a variety of LLMs, demonstrating that IHs are crucial components for many reasoning tasks. Further, we propose that subspace matching—and more broadly common bridge representation hypothesis—as the compositional mechanism.
@@
Historically, studies of generalization on novel domains focus on extrapolation, distribution shift, domain adaptation, etc. Since GPT-3 [16], recent studies of OOD generalization focus on arithmetic and algebraic tasks [41, 90], formal language and deductive reasoning [66, 62, 71, 79], learning boolean functions [1], etc. Our copying task is an example of length generalization [6, 92]. Compositional tasks are strongly related to reasoning abilities of LLMs, such as arithmetic tasks [23, 40], formal language [31], logical rules [14, 87], and so on. Despite recent insights [8, 23, 81, 18], there is a lack of systematic and quantitative analysis. Our work is a first step toward understanding compositional capabilities of LLMs in a systematic way.
@@
Linear Representation Hypothesis (LRH) states that monosemantic (meaning basic or atomic) concepts are represented by linear subspaces (or half-spaces) in embeddings [9, 24, 57]. This hypothesis is empirically verified not only in simpler statistical models but also in LLMs [49, 58]. The LRH treats subspaces as fundamental units for representing linguistic concepts, but it does explain how models solve reasoning tasks or achieve OOD generalization. Our CBR hypothesis furthers this view by linking intermediate outputs in compositions to interpretable subspaces.
@@ 6 Limitations and future work
Second, we did not develop insights to explain the emergence of the bridge subspace during training. The sharp transition in prediction accuracy is related to the emergent abilities of LLMs observed in broader contexts [82]. In simpler settings, feedforward networks learning algebraic rules also exhibit phase transitions from memorization to generalization, a phenomenon known as Grokking [60, 52, 91, 45, 44, 47]. Analyzing the training dynamics of Transformers for copying and other compositional tasks would be an interesting avenue for further research.
@@ B.3 One-layer Transformer
Figure 1 shows that one-layer Transformers do not learn the rule of copying under 20K training steps. We used the same model architecture and hyperparameters for training as detailed in Section D.1.
@@ B.4 Larger Transformers
The results of the three models are similar to the minimal model (2-layer 1-head Transformer with no MLP) we analyzed in the paper. See plots in Section C.1.
@@ B.5 Details about models
There are two major variants of positional embedding: absolute positional embedding (APE) and rotary embedding (RoPE). APE is proposed in the original Transformer paper [76] and used widely in early LLMs such as GPT-2. The input vector to a Transformer is obtained by adding a token-specific embedding vector and a position-specific embedding vector, thus encoding both token information and positional information. RoPE [70] is often used in recent LLMs such as Llama-2. The positional information is not encoded at the input layer; instead, RoPE directly modifies the attention by replacing ${\bm{W}}^{q}({\bm{W}}^{k})^{\top}$ in Eq. 8 with ${\bm{W}}^{q}{\bm{R}}_{\Delta t}({\bm{W}}^{k})^{\top}$ where ${\bm{R}}_{\Delta t}\in\mathbb{R}^{d_{\mathrm{head}}\times d_{\mathrm{head}}}$ is a matrix that depends on relative distance of two tokens. In particular, if $\Delta t=0$ then ${\bm{R}}_{\Delta t}={\bm{I}}_{d}$. Our analysis in Section 3 and 4 essentially treats $\Delta t=0$.
Section C.5 explores the impact of ${\bm{R}}_{\Delta t}$ on our results. We did not find significant differences when we repeat our measurements with a nonzero $\Delta t$.
@@ B.6 Details about measurements
Appendix C Additional results for the synthetic example
C.1 Larger models show similar sharp transitions
We believe that the 2-layer 1-head Transformer with no MLP is a minimal yet representative model for analyzing the sharp transition. We show similar results on three larger Transformer models in Figure 9 to 12.
Figure 12: Progress measures for 2-layer 1-head Transformer with MLP.
C.2 Progress measures under varying pool sizes
We consider the same model/training settings as in Section 3 but we choose two different pool sizes $S=1000$ and $S=100$. In Figure 13 and 14, the progress measures under $S=1000$ are similar to Figure 2 but very different under $S=100$.
C.3 Attention matrices
The heatmaps in Figure 3 indicate that the memorizing model ($S=100$) are qualitatively different from that generalizing model ($S=1000$). In Figure 15 and 16, we further plot attention matrices ${\bm{A}}$ based on one OOD sequence at training step 16K to support our claim that the memorizing model does not learn compositional structure for OOD generalization.
C.4 Measurements for subspace matching
$\sigma_{\max}({\bm{U}}^{\top}{\bm{V}})$
  • Vary the rank parameter $r$.
We find that the two alternative measurement yield qualitatively similar results; see Figure 17 and 18.
C.5 Rotary positional embedding
${\bm{W}}_{\mathrm{QK},\Delta t}={\bm{W}}^{q}{\bm{R}}_{\Delta t}({\bm{W}}^{k})^{\top}$
C.6 Connection to spiked matrices
Appendix D Model and task details for experiments with LLMs
D.1 Models
@@
  • Llama2-70B: 80 layers, 64 attention heads, and a hidden size of 8192.
  • Llama3-70B: 80 layers, 64 attention heads, and a hidden size of 8192.
@@
D.2 Tasks
We provide details including setup, measurement and prompt examples for tasks introduced in our main paper: Fuzzy Copying, Indirect Object Identification (IOI), In-Context Learning (ICL), GSM8K, and GSM8K-rand.
Fuzzy copying
The task is evaluated on 100 samples, each containing 10 in-context examples. For each sample, the result is determined by whether the model correctly predicts the next two words. Two metrics are used for evaluation in IOI: accuracy and probability. For accuracy, each sample is evaluated based on whether the model correctly predict all the tokens in the answer. The mean accuracy across all samples is reported. For probability, we calculate the cumulative probability of the model predicting the correct tokens for each sample. The mean probability across all samples is then reported.
Here is an example prompt for the Fuzzy Copying task.
bear snake fox poppy plate butterfly caterpillar boy
@@ D.2 Tasks
IOI
The task is evaluated on 100 samples. Two metrics are used for evaluation in IOI: accuracy and probability. For accuracy, each sample is evaluated based on whether the model assigns a higher probability to the correct name among the two options. The mean accuracy across all samples is reported. For probability, we calculate the cumulative probability of the model predicting the correct tokens for each sample. The mean probability across all samples is then reported.
Here are example prompts for the IOI task, presented in both the English and symbolized versions.
@@
Then, Anna and Matthew went to the station. Anna gave a basketball to
@@
Then, &^ and #\$ went to the station. &^ gave a basketball to
ICL
The task is evaluated on 100 samples, each containing 20 in-context examples. For each sample, the result is determined by whether the model correctly predicts the next category. Two metrics are used for evaluation in IOI: accuracy and probability. For accuracy, each sample is evaluated based on whether the model correctly predict all the tokens in the answer. The mean accuracy across all samples is reported. For probability, we calculate the cumulative probability of the model predicting the correct tokens for each sample. The mean probability across all samples is then reported.
Here are example prompts for the ICL task, presented in both the English and symbolized versions.
@@
baseball is sport, celery is plant, sheep is animal,
@@
baseball is \$#, celery is !%, sheep is &*, volleyball is \$#,
@@
GSM8K
The task is evaluated on 100 samples, each containing 10 CoT examples of GSM questions. For each sample, the answer is extracted using regular expression matching for the pattern #### {solution}, and the result is determined by whether the model’s solution exactly matches the extracted answer. We report the mean accuracy across all samples. Below is an example prompt. For demonstration purposes, we include only two in-context examples.
Question: Jack has a stack of books that is 12 inches thick. He knows from experience that 80 pages is one inch thick. If he has 6 books, how many pages is each one on average?
Answer: There are 960 pages because 80 x 12 = <<80*12=960>>960
Each book is 160 pages because 960 / 6 = <<960/6=160>>160
#### 160
Question: Joseph invested \$1000 into a hedge fund. The fund promised a yearly interest rate of 10%. If he deposited an additional \$100 every month into the account to add to his initial investment of \$1000, how much money will he have in the fund after two years?
Answer: For the first year, Joseph will have invested \$1000 + (\$100 * 12) = \$<<1000+100*12=2200>>2200.
The interest calculated for the first year will be \$2200 * 10% = \$<<2200*10*.01=220>>220.
The total value of the investment for the first year will be \$2200 + \$220 = \$<<2200+220=2420>>2420.
For the second year, the total invested will be \$2420 + (\$100 * 12) = \$<<2420+100*12=3620>>3620.
The interest calculated after the second year will be \$3620 * 10% = \$<<3620*10*.01=362>>362.
Therefore, Josephs investment in the mutual fund will be worth \$3620 + \$362 = \$<<3620+362=3982>>3982.
#### 3982
<<<<< EIGHT MORE IN-CONTEXT QUESTION-ANSWER PAIRS >>>>>
Question: Craig has 2 twenty dollar bills. He buys six squirt guns for \$2 each. He also buys 3 packs of water balloons for \$3 each. How much money does he have left?
GSM8K-rand
To construct the GSM8K-rand dataset, we sample 12 questions from GSM8K and use them to create 12 templates. Specifically, for each question, a template is generated by replacing numbers and names in the problem statement and CoT deduction process with references. These references are drawn from a pool of 40 symbols, making the problems unlikely to appear in the training data. Additionally, the sampled questions are deliberately chosen from those correctly predicted by the Llama2-7B model. This selection helps establish a higher baseline and allows for a clearer observation of accuracy reduction during the intervention.
Regarding the accuracy evaluation, each prompt consists of 10 unique template realizations followed by one question. All other details remain the same as those used for GSM8K. Below is an example prompt. For demonstration purposes, we include only two in-context examples.
Question: A boy has 15 #\$. His brother has 2 fewer #\$ than he has. How many #\$ do they have together?
Answer:
His brother has 15 - 2 = <<15-2=13>>13 #\$.
Together, they have 15 + 13 = <<15+13=28>>28 #\$.
#### 28
Question: There is space for 12 &! in the box. If there are 6 &! missing from the box, how many pairs of &! are in the box?
Answer:
In the box there are 12 &! - 6 &! = <<12-6=6>>6 &!.
Dividing into pairs we have 6 &! / 2 &!/pair = 3 pairs of &!
#### 3
<<<<< EIGHT MORE IN-CONTEXT QUESTION-ANSWER PAIRS >>>>>
Question: &* is cutting up wood for his wood-burning stove. Each \$@ tree makes 90 logs, each @\$ tree makes 63 logs, and each @^ tree makes 70 logs. If &* cuts up 7 \$@ trees, 4 @\$ trees, and 4 @^ trees, how many logs does he get?
Appendix E Experiment details for the Common bridge representation hypothesis
E.1 Details about PTHs and IHs in LLMs